OpenStudy (anonymous):

Someone please help!! What are the coordinates of the foci of the conic section shown below? (y+2)^2/25- (x-3)^2/4=1 A. (3±√21, -2) B. (3, -2±√29) C. (3, -2±√21) D. (3±√29, -2)

OpenStudy (anonymous):

C. (3, -2±√21)

OpenStudy (anonymous):

thank you!!

OpenStudy (anonymous):

:)

OpenStudy (stacey):

a=5 and b=2 c^2 = 25 + 4 = 29 So we add and subtract √29 to the y-coordinate.

OpenStudy (stacey):

The center is (3, -2) The foci are at (3, -2±√29).